All of number 3 please including letters d and e at the bottom
(a) Use the method of moments to find a point estimate for p. 3.(100] 6.5-3. The midterm and final exam scores of 10 students in a statistics course are tabulated as shown. (a) Calculate the least squares regression line for these data. (b) Plot the points and the least squares regression line on the same graph. (c) Find the value of σ2. Midterm Final Midterm Final 70 87 67 70 64 74 73 83 79 91 94 74 79 84 98 80 96 (d) Compute and plot residuals (e) What is the expected final exam score of a student whose midterm score was x = 68?
Answer #1

a) The following data are provided:

Midterm Final
70 87
74 79
80 88
84 98
80 96
67 73
70 83
64 79
74 91
82 94

The independent variable is Midterm, and the dependent variable is Final. In order to compute the regression coefficients, the following table needs to be used:

Midterm Final Midterm*Final Midterm2 Final2
70 87 6090 4900 7569
74 79 5846 5476 6241
80 88 7040 6400 7744
84 98 8232 7056 9604
80 96 7680 6400 9216
67 73 4891 4489 5329
70 83 5810 4900 6889
64 79 5056 4096 6241
74 91 6734 5476 8281
82 94 7708 6724 8836
Sum = 745 868 65087 55917 75950

Based on the above table, the following is calculated:

745 74.5 10 に!

868 86.8

7n. ssx x 「X2-1 X, | = 55917-7452/10 414.5

SSYY=Σ12_n(ΣΥ 〉 Yİ | = 75950-8682/10 607.599

Y; | 65087-745 × 868/ 10 421

Therefore, based on the above calculations, the regression coefficients (the slope m, and the y-intercept n) are obtained as follows:

SSxY 421 SSrx414.5 414.5 1.0157

n-y-X . m = 86.8-74.5 × 1.0157 11.1317

Therefore, we find that the regression equation is:

Final = 11.1317 + 1.0157 * Midterm

y = 11.1317 + 1.0157 * x

b)

Graphically:

Scatter Plot and Regression Line 103 98 93 LL 83 78 73 680 62 64 66 68 70 72 74 7678 80 82 84 8688 Midtern -Regression equati

c)

을 607.59-44.2249850875 Syy

d)

Midterm(x) Final(y) y^ = 11.1317 + 1.0157 * x residual = y - y^
70 87 82.2307 4.7693
74 79 86.2935 -7.2935
80 88 92.3877 -4.3877
84 98 96.4505 1.5495
80 96 92.3877 3.6123
67 73 79.1836 -6.1836
70 83 82.2307 0.7693
64 79 76.1365 2.8635
74 91 86.2935 4.7065
82 94 94.4191 -0.4191
Sum = 745 868 868.0135 -0.0135

1100 20 30 40 50 60 70 80 90 Residual Plot

e)

When Midterm = x = 68

Final = y = 11.1317 + 1.0157 * 68 = 80.1993

Please upvote! Thanks!